# The Tautochrone of Huygens and Abel: From Constructive Geometry to Fractional Calculus

### v1

@@ Abstract
Keywords: Isoperimetric problems; variational calculus; cycloid; integral equations; fractional calculus.
MSC: 01-01, 01-02, 45E10.
1 Introduction
Problems involving extremes (maximums and minimums) are old and persistent. The very use of terms like “law of least effort”, “path of least resistance”, “shortest distance between two points”, etc., is part of our common sense of economy or simplicity in the actions of nature [1]. The first problem of finding an extreme that is known to us is associated with the name of the Phoenician Princess Dido, queen of Carthage — immortalized by the Latin poet Virgil in the Aeneid, and later by the English composer Henry Purcell (1659–1695) in the musical drama Dido & Aeneas. The legend says that the princess, having emigrated to North Africa after being defrauded of her belongings by her brother, Pygmalion, was promised by Hiarbas, the native chief of the region, as much land as she could enclose in a cowhide. Cleverly, she cut the cowhide into a very long strip and placed it between herself and the sea, creating a territory large enough to build the city of Carthage [2]. The typical solution found by Dido is to arrange the strip in a circular shape. It is, therefore, an isoperimetric problem, one of many others that will be encountered throughout history. Another problem of this nature is associated with the name of Heron of Alexandria (10–70 A.D.). He is credited with deriving the law of reflection from the principle that the light ray emitted from point A reaches point B after being reflected by a mirror, following the path that takes the least time. This principle was extended by Fermat to derive Snell–Descartes’ law of refraction [1].
@@ 1 Introduction
$\psi(a)=\int_{x=0}^{x=a}\frac{f(x)}{\sqrt{a-x}}\,dx,$ (1)
@@
Here, we aim to highlight the scientific contributions of these two mathematicians, using the tautochrone problem as a guiding thread. Some biographical details are included to underscore the contrast between their backgrounds, beginning in childhood, when both had already shown a rare and precocious talent for mathematical science and abstract thinking. By focusing on how each of them approached the problem — separated by two centuries — we also trace how analysis became increasingly sophisticated in addressing important and advanced problems in physics and mathematics. Special attention is given, in the final part of the paper, to the manners in which Abel’s work paved the way for the development of key concepts in fractional calculus — matured in the years that followed [3].
2 The Geometric Construction of Huygens
@@
2.1 The Tautochronism of the Cycloid
Because astronomical investigations require increasingly accurate time measurements, Huygens devoted considerable effort to the problem. In 1656, he patented the first pendulum clock, significantly improving the precision of timekeeping. This innovation was later described in his 1658 publication Horologium (“The Clock”). Although his investigations into centrifugal force also date back to this period (1659), it was in his major work of 1673, Horologium Oscillatorium sive de motu pendulorum (“The Pendulum Clock, or On the Motion of Pendulums”), that he developed a full theoretical account of pendular motion. It is also in this treatise that the modern formula for the centrifugal force in uniform circular motion first appears. The Horologium Oscillatorium is often regarded as one of the most important work on mechanics prior to Newton’s Principia (1687). In the book, Huygens derived the formula for the period of a simple pendulum under the small-angle approximation — a classical result encountered in the Galileo’s work. Huygens went further by investigating the precise curve along which a pendulum would swing with a period that is truly independent of its amplitude. His solution identified this path as a cycloid.
This remarkable property is known as the tautochronism of the cycloid — from the Greek words ταυτό (equal, same) and ϰρόνος (time) — and it requires that the pendulum follow an inverted cycloidal trajectory. The tautochrone problem is addressed in Part II, specifically in Propositions XVI through XXVI. Let us now proceed by sketching some of the representative steps of his derivation, focusing the last two propositions, starting with Proposition XXV. It states that [4]:
Proposition XXV
In a cycloid with a vertical axis, and with the vertex seen to be the lowest point, the times of descent for some body, on leaving any point on the cycloid from rest until it reaches the lowest point at the vertex, are equal to each other; and this time has the same ratio to the time of fall along the whole axis of the cycloid as the semi-circumference of the circle to the diameter.
To reproduce his construction, we need the help of Fig. 1, part of the Huygens’ book [5].
@@ 2.1 The Tautochronism of the Cycloid
For a better understanding of the discussion that follows, here is a summary of the elements that appear in Fig. 1 and their meaning:
@@
The cycloid shall be ABC the vertex of which is seen to be A pointing downwards, with the vertical axis truly AD, and by taking some point on the cycloid, such as B, a body is free to fall from natural forces along the arc BA, or along a surface thus curved into this shape. Huygens says that the time of descent of this body to be as to the time of fall along the axis DA, thus as the semi-circumference of the circle to the diameter. From which demonstration, the time of descent along any arc of the cycloid to finish at A also, are agreed to be equal to each other [6]. Huygens then states that the descent time of this body is, in relation to the fall time along the DA axis, as the semicircle is in relation to the diameter. This proportion stated by Huygens — between the descent time on the cycloid and the vertical fall time along the axis — is the expression of tautochronism: in fact, the proportionality constant is $\pi/2$, since the perimeter of the semicircle is $\pi\,r$, while its diameter is $2r$.
The discussion continues by diving into the details of the geometric construction. Huygens explains that a semicircle is drawn on the axis DA. This semicircle intersects the line BF at point E, which is parallel to the base DC. The line EA is then drawn, and the line BG is drawn parallel to it, which turns out to be a tangent to the cycloid at point B. In fact, this line also intersects the line drawn horizontally from point A at point G. Additionally, a semicircle FHA is drawn on the line FA. With this setup, Huygens states that the time it takes to descend along the arc of the cycloid BA is to the time of uniform motion along the line BG (with half the speed of BG) as the arc of the semicircle FHA is to the straight line FA. In other words, the time for uniform motion along BG is the same as the time it would take to descend along the same line BG, or along EA (which is parallel and equal to BG), or equivalently, the time to fall under gravity along the axis DA, as shown in Proposition 6 by Galileo [7]. Therefore, the time along the arc BA is related to the time of descent along the axis DA in the same way the circumference of the semicircle FHA is related to the diameter FA.
The last proposition dedicated to this problem also refers to Fig. 1 and is stated as follows [8]:
@@
The Huygens’ explanation of the construction may be paraphrased as follows [9]. Consider the straight line HI intersecting the tangent BG at point K, and the axis DA at point L. The time it takes to travel along the arc BA is compared to the time it takes for uniform motion along BG with half the speed of BG. This comparison is similar to how the arc FHA compares to the straight line FA. The time of uniform motion along BG is compared to the time of uniform motion along BK, with the same half-speed of BG. The ratio of these times is the same as the ratio of BG to BK, which is equivalent to the ratio of FA to FL. The time for uniform motion with the same speed along BK is compared to the time it takes to move along the arc BI, which is similar to how the length FL compares to the arc FH.
As a result, the time to travel along the arc BA is equal to the time to travel along the arc BI, and this is in the same proportion as the arc FHA is to the arc FH. Finally, by dividing and converting, the ratio of the time along BI to the time along IA after BI is the same as the ratio of the arc FH to HA. We underline that both propositions end by the Q.E.D. statement in the original work because for Huygens the geometrical constructions he employs are rigorous method to demonstrate the validity of a physical result. This is indeed the case as we can show hereafter by recalling the problem using simpler language and more updated mathematical notation.
To proceed this way, let’s first remember that Huygens started with the geometric generation of the cycloid, i.e., he considered a circle of radius $r$ rolling on a horizontal surface without slipping. A fixed point on the edge of the circle traces a path that defines the cycloid. The resulting curve is symmetric, with its lowest point at the vertex. When the curve is inverted (with the concavity facing upwards), it gives the ideal profile of a tautochrone pendulum. Huygens’ analysis leads us to think that, for the descent time along a curve to be independent of the starting point, the curve must have the property of equating the acceleration of a body along it with the appropriate free fall acceleration for comparison purposes.
2.2 The Tautochrone as a Variational Problem
Using a modern approach, we could say that he wants to find a curve such that
$\int\frac{ds}{\sqrt{2gy}}=\textrm{constant},$ (2)
where $ds$ is an element of the curve, $g$ is the acceleration due to gravity, and $y$ denotes the vertical axis in a Cartesian system. In this approach, the deduction is based on comparing two motions, namely, the motion along the inverted cycloid (frictionless sliding) and uniform circular motion, i.e., motion with constant speed along a circle. Based on these premises, Huygens establishes a correspondence between the vertical height of the particle on the cycloid and the angular position of a point moving uniformly along a circle. He then shows that, if one considers an auxiliary circle that generates the cycloid (by rolling), the time it takes for a point to reach a certain height on the cycloid coincides with the time a point in uniform motion takes to reach the same angle on the circle — provided the variables are correctly associated. In a few words, the projection of uniform circular motion along the generatrix of the cycloid coincides with the free fall motion along the cycloid — and since uniform circular motion is isochronous, the motion along the cycloid will also be isochronous!
In proceeding, Huygens notes that a common pendulum (one suspended by an ideal string) is not isochronous: its period depends on the amplitude of the oscillation. To resolve this, he proposes restricting the motion of the string so that the pendulum, when oscillating, follows an inverted cycloid rather than a circular arc. This ensures that, regardless of the amplitude, the oscillation period of the pendulum remains exactly constant. With these ingredients, Huygens then demonstrates that the inverted cycloid is the only curve for which the descent time is independent of the starting point. In fact, the descent time from any point to the vertex is
$T=\pi\sqrt{\frac{r}{g}},$ (3)
which coincides with the period of a cycloidal pendulum. Finally, the ratio between this time and the free fall time along the cycloid’s axis is given by
$\frac{T_{\textrm{cycloid}}}{T_{\textrm{free \, fall}}}=\frac{\pi}{2}.$ (4)
In other words and somewhat more formal terms, we wish to find the curve $y(x)$ such that
$\int_{A}^{B}\frac{ds}{v(y)}=\textrm{constant},$ (5)
where $v(y)$ is the velocity of the particle along $y$. In this case, we have
$ds=\sqrt{1+\left(\frac{dy}{dx}\right)^{2}}dx=\sqrt{1+y^{\prime 2}}dx$ (6)
and, also, the velocity is given by $v(y)=\sqrt{2gy}$, a result derived using the conservation of the mechanical energy. Thus, the descent time for the particle can be defined as
$T=\int\frac{\sqrt{1+y^{\prime 2}}}{\sqrt{2gy}}dx,$ (7)
which is the same integral that appears in the brachistochrone problem — the curve of minimum time between two points.
The starting point of this investigation, in fact, was a challenge posed by Johann Bernoulli in 1696. It consisted of finding the trajectory followed by a body starting from rest at a given point, in a gravitational field, along a frictionless curve, until the final point. The curve is known as the brachistochrone, from Greek: βράϰιςτος (shortest) and ϰρόνος (time). The problem was tackled (and solved) by Jakob I Bernoulli, in an article dedicated to solving one of the problems proposed by his brother: Solutio problematum fraternorum (Solution to the brother’s problem), which he titled De curva celerrimi descensus (literally: the curve of fastest descent) [10]. Euler later solved the same problem, thereby founding the calculus of variations [11, 12]. Later, Lagrange revolutionized the field by introducing the $\delta$ algorithm and incorporated variational methods into a much more general mathematical framework [13, 14].
As can be seen from the excerpts of Huygens’ demonstration, none of these equations were employed in establishing his results. We have introduced them here merely to bring Huygens’ language closer to the one we use in the study of the fundamental problems of mechanics. This fact, by itself, illustrates the importance that analytical methods have played in mechanics. Indeed, in the period between the publication of Newton’s Principia in 1687 and the emergence of Lagrange’s Méchanique Analytique in 1788, a highly elaborate language in mathematical physics developed — an outcome of the effort to integrate mechanics into mathematical analysis. It was a fertile period during which scientists of great stature, such as Leibniz, Varignon, Jacob I and Johann I Bernoulli (and his son Daniel), Euler, d’Alembert, Laplace and Lagrange, among others, devoted themselves wholeheartedly to this endeavor. The work of these men led to the emergence of analytical mechanics, practically in the form we still know and that is presented in our textbooks today.
3 The Integral Equations of Abel
@@
Niels Henrik Abel was a Norwegian mathematician renowned for his groundbreaking work in algebra. He is best known for proving that the general equation of the fifth degree (quintic equations) cannot be solved by radicals, a result now known as Abel’s impossibility theorem (Abel–Ruffini theorem) [15, 16]. Abel’s work laid the foundation for much of modern algebra and group theory. Despite dying young at 26, Abel made significant contributions to mathematics, including studies on elliptic functions and the development of Abelian integrals. His legacy is celebrated in the mathematical community, and the prestigious Abel Prize is named in his honor.
@@
In 1823, Abel solved the tautochrone problem by using a fractional integration of order $\alpha=1/2$ [17] and, in 1826, extended the approach to $\alpha\in(0,1)$ [18]. In his 1826 work, Abel presents a different solution method, based on the use of the properties of Euler’s gamma function. As emphasized in Ref.[20], it is a more concise and elegant method which, however, abandons the path opened toward fractional calculus, characteristic of his 1823 paper. Indeed, as we briefly discuss now, the solution proposed by him to the tautochrone problem involves the pioneering use of an integral operator of arbitrary order, and is the first explicit solution known of an integral equation[19].
3.1 Abel’s Solution: Remarkable Theorem
On the first page of Abel’s 1823 article, dedicated to definite integrals, we read [17]:

### v2

@@ Abstract
Keywords: Isoperimetric problems; variational calculus; integral equations; fractional calculus
I Introduction
Problems involving extremes (maximums and minimums) are old and persistent. The very use of terms like “law of least effort”, “path of least resistance”, “shortest distance between two points”, etc., is part of our common sense of economy or simplicity in the actions of nature Fermat . The first problem of finding an extreme that is known to us is associated with the name of the Phoenician Princess Dido, queen of Carthage — immortalized by the Latin poet Virgil in the Aeneid, and later by the English composer Henry Purcell (1659–1695) in the musical drama Dido & Aeneas. The legend says that the princess, having emigrated to North Africa after being defrauded of her belongings by her brother, Pygmalion, was promised by Hiarbas, the native chief of the region, as much land as she could enclose in a cowhide. Cleverly, she cut the cowhide into a very long strip and placed it between herself and the sea, creating a territory large enough to build the city of Carthage Kelvin . The typical solution found by Dido is to arrange the strip in a circular shape. It is, therefore, an isoperimetric problem, one of many others that will be encountered throughout history. Another problem of this nature is associated with the name of Heron of Alexandria (10–70 A.D.). He is credited with deriving the law of reflection from the principle that the light ray emitted from point A reaches point B after being reflected by a mirror, following the path that takes the least time. This principle was extended by Fermat to derive Snell–Descartes’ law of refraction Fermat .
@@ I Introduction
$\psi(a)=\int_{x=0}^{x=a}\frac{f(x)}{\sqrt{a-x}}\mskip 3.0mudx,$ (1)
@@
Here, we aim to highlight the scientific contributions of these two mathematicians, using the tautochrone problem as a guiding thread. Some biographical details are included to underscore the contrast between their backgrounds, beginning in childhood, when both had already shown a rare and precocious talent for mathematical science and abstract thinking. By focusing on how each of them approached the problem — separated by two centuries — we also trace how analysis became increasingly sophisticated in addressing important and advanced problems in physics and mathematics. Special attention is given, in the final part of the paper, to the manners in which Abel’s work paved the way for the development of key concepts in fractional calculus — matured in the years that followed Ross1977 .
II The Geometric Construction of Huygens
@@
II.1 The Tautochronism of the Cycloid
Because astronomical investigations require increasingly accurate time measurements, Huygens devoted considerable effort to the problem. In 1656, he patented the first pendulum clock, significantly improving the precision of timekeeping. This innovation was later described in his 1658 publication Horologium (“The Clock”). Although his investigations into centrifugal force also date back to this period (1659), it was in his major work of 1673, Horologium Oscillatorium sive de motu pendulorum (“The Pendulum Clock, or On the Motion of Pendulums”), that he developed a full theoretical account of pendular motion Huygens1673 . It is also in this treatise that the modern formula for the centrifugal force in uniform circular motion first appears. The Horologium Oscillatorium is often regarded as one of the most important work on mechanics prior to Newton’s Principia (1687). In the book, Huygens derived the formula for the period of a simple pendulum under the small-angle approximation — a classical result encountered in the Galileo’s work. Huygens went further by investigating the precise curve along which a pendulum would swing with a period that is truly independent of its amplitude. His solution identified this path as a cycloid. In a few words, we can say that Huygens showed the cycloid is a tautochrone in a uniform gravitational field.
@@ II.1 The Tautochronism of the Cycloid
1
Indeed, this remarkable property is known as the tautochronism of the cycloid — from the Greek words ταυτό (equal, same) and ϰρόνος (time) — and it requires that the pendulum follow an inverted cycloidal trajectory. The tautochrone problem is addressed in Part II, specifically in Propositions XVI through XXVI. To reproduce his construction, we need the help of Fig. 1, part of the Huygens’ book Huygens1673b . For a better understanding of the discussion that follows, here is a summary of the elements that appear in Fig. 1 and their meaning:
@@
Let us now proceed by sketching some of the representative steps of his derivation, focusing the last three propositions, starting with a brief summary of the Proposition XXIV, from which follow almost immediately the Propositions XXV and XXVI. This is useful because Huygens’ solution rests on a specific geometric property of the cycloid, established in Proposition XXIV, which he then applies in the culminating Propositions XXV and XXVI. In this perspective, Proposition XXIV is the geometric foundation of his derivation because, in this framework, he first proves a critical lemma: for any point on an inverted cycloid, the ratio of the arc length from that point to the vertex to the straight chord connecting them is constant. More precisely, for any point I on the cycloid arc BA (see Fig. 1),
$\frac{\text{arc}IA}{\text{chord }IA}=\frac{\text{semicircumference of the generating circle}}{\text{its diameter}}=\frac{\pi}{2}.$
This constant arc–to–chord ratio is a kind of kinematic signature of the cycloid. It is not an obvious identity; Huygens derives it through an intricate geometric construction involving the cycloid’s evolute Huygens1673b . This proposition provides the precise geometric tool needed to translate distances into times.
Using the geometric identity established in Proposition XXIV, Huygens proves the tautochrone property in Proposition XXV:
Proposition XXV In a cycloid with a vertical axis, and with the vertex seen to be the lowest point, the times of descent for some body, on leaving any point on the cycloid from rest until it reaches the lowest point at the vertex, are equal to each other; and this time has the same ratio to the time of fall along the whole axis of the cycloid as the semi-circumference of the circle to the diameter.
In the accompanying construction (Fig. 1), Huygens demonstrates that the descent time along the cycloid arc BA is proportional to the arc length of the auxiliary semicircle FHA, while the reference free-fall time along the axis DA is proportional to its diameter FA Bruce . The ratio between these times—the constant $\pi/2$ from Proposition XXIV—thus quantifies the tautochrone period: it equals $(\pi/2)$ multiplied by the free-fall time along the axis.
The kinematic reasoning proceeds as follows: the time along arc BA relates to the time of uniform motion along line BG (with half the speed acquired in free fall) as arc FHA relates to line FA. By Galileo’s laws Galileo , that uniform motion time equals the free-fall time along DA. Consequently, the cycloidal descent time is permanently tied to the $\pi/2$ ratio of the generating circle.
Huygens then examines the internal distribution of this constant time in Proposition XXVI Bruce :
@@
This proposition reveals the geometric clockwork inside the isochronous motion: time elapsed maps linearly onto the arc of the generating circle. For any intermediate point I, the ratio of times (BI to IA) equals the ratio of circular arcs (FH to HA). In modern terms, this expresses the differential condition that ensures the total time remains constant; each segment of the descent contributes a duration proportional to a corresponding arc segment on the auxiliary circle.
Thus, Huygens’s synthetic method follows a clear architecture:
  • Proposition XXIV provides the fundamental geometric identity (arc/chord = $\pi/2$).
  • Proposition XXV uses this identity to prove and quantify global isochrony.
  • Proposition XXVI derives the internal temporal structure, confirming the self-consistency of the geometric clock.
Crucially, Huygens did not derive the cycloid from the tautochrone condition. Instead, starting from the known curve (generated by a rolling circle of radius $r$ and then inverted), he verified its isochrony through geometric proportions and kinematic equivalences. His reasoning was strictly comparative—relying on Galileo’s laws and the fixed ratio
$\frac{\text{arc IA}}{\text{chord IA}}=\frac{\pi}{2},$
and never invoked differential equations or variational principles. What a modern reader might interpret as the curve’s curvature adjusting tangential acceleration is, for Huygens, the consequence of a geometric proportion between arcs, chords, and times.
II.2 From Huygens’s Geometry to Modern Analytic Language
Huygens’s 1673 demonstration of the tautochrone property was purely geometric. As we have seen, he established that for an inverted cycloid, the descent time from any point to the vertex equals the constant $T=\pi\sqrt{r/g}$, where $r$ is the radius of the generating circle. His proof relied entirely on proportions between arc lengths, chords, and times—never on analytic expressions.
At the heart of Huygens’s reasoning is a kinematic correspondence: the descent along the cycloid arc BA maps linearly onto uniform motion along the arc of its generating circle. Specifically, he shows that the vertical height of the particle on the cycloid corresponds to the angular position of a point moving uniformly along the auxiliary circle. Since uniform circular motion is isochronous, the descent along the cycloid inherits the same property. This geometric insight also explains why constraining an ordinary pendulum with “cycloidal cheeks” renders its period amplitude-independent Huygens1673 ; Bruce .
In modern analytic terms, the tautochrone condition can be expressed as follows. For a particle sliding without friction from rest at height $y_{0}$, conservation of energy gives $v=\sqrt{2g(y_{0}-y)}$. The descent time is therefore
$T=\int_{0}^{y_{0}}\frac{ds}{\sqrt{2g(y_{0}-y)}}=\text{constant},$ (2)
independent of $y_{0}$. With $ds=\sqrt{1+(dy/dx)^{2}}\mskip 3.0mudx$, this becomes
$T=\int\frac{\sqrt{1+y^{\prime 2}}}{\sqrt{2g(y_{0}-y)}}\mskip 3.0mudx,$ (3)
taken along the unknown curve $y(x)$.
This formulation reveals the tautochrone as an integral constraint to be inverted—exactly the approach Abel would take in 1823 Abel1823 . For Huygens, however, the cycloid was not derived from this equation but verified against it through geometric means. His result can be summarized as
$\frac{T_{\text{cycloid}}}{T_{\text{free fall}}}=\frac{\pi}{2},$
a ratio that emerged from the geometry of the cycloid and its generating circle, not from manipulating integrals.
Connection to the Brachistochrone and Calculus of Variations.
A different but related problem—the brachistochrone (from Greek βρἑϰιςτος, shortest, and ϰηρῗνος, time)—was posed by Johann Bernoulli in 1696, two decades after Huygens’s work Bernoulli . Bernoulli asked for the curve of fastest descent between two given points (not equal-time descent from any starting point). The brachistochrone functional,
$T_{\text{brach}}=\int\frac{\sqrt{1+y^{\prime 2}}}{\sqrt{2gy}}\mskip 3.0mudx,$
mathematically resembles the tautochrone integral, though the problems are distinct: the brachistochrone minimizes time between two fixed points, while the tautochrone demands time invariance for variable starting points. Remarkably, both yield the same curve—the cycloid.
Jakob Bernoulli’s 1697 solution to the brachistochrone Bernoulli , followed by Euler’s systematization in 1744 Euler1744 ; Fraser1994 and Lagrange’s $\delta$-algorithm in the 1760s Lagrange-LF ; Fraser-LF , laid the foundations of the calculus of variations. This analytic development stands in stark contrast to Huygens’s geometric synthesis, highlighting the profound transformation of mechanics between 1673 and 1823—from synthetic verification to analytic derivation.
III The Integral Equations of Abel
@@
Niels Henrik Abel was a Norwegian mathematician renowned for his groundbreaking work in algebra. He is best known for proving that the general equation of the fifth degree (quintic equations) cannot be solved by radicals, a result now known as Abel’s impossibility theorem (Abel–Ruffini theorem) Abel1824 ; Abel1826b . Abel’s work laid the foundation for much of modern algebra and group theory. Despite dying young at 26, Abel made significant contributions to mathematics, including studies on elliptic functions and the development of Abelian integrals. His legacy is celebrated in the mathematical community, and the prestigious Abel Prize is named in his honor.
@@
In 1823, Abel solved the tautochrone problem by using a fractional integration of order $\alpha=1/2$ Abel1823 and, in 1826, extended the approach to $\alpha\in(0,1)$ Abel1826 . In his 1826 work, Abel presents a different solution method, based on the use of the properties of Euler’s gamma function. As emphasized in Ref. Podlubny-Magin , it is a more concise and elegant method which, however, abandons the path opened toward fractional calculus, characteristic of his 1823 paper. Indeed, as we briefly discuss now, the solution proposed by him to the tautochrone problem involves the pioneering use of an integral operator of arbitrary order, and is the first explicit solution known of an integral equation Lutzen1982 .
III.1 Abel’s Solution: A Remarkable Theorem
Abel opens his 1823 paper with a statement that highlights his methodological shift Abel1823 :